# LEKA: LLM-Enhanced Knowledge Augmentation

### v1

@@ Introduction
Experiment Settings
Datasets and Domains. We evaluate our method primarily on four datasets of medical and economic domains: (1) Breast Cancer Wisconsin (Diagnostic) (BCW) (Wolberg et al. 1995), (2) Heart Disease (HD) (Janosi et al. 1988), (3) Vehicle Insurance Data (VID) (Bhatt 2019), and (4) Telco Customer Churn (TCC) (Zhang, Moro, and Ramos 2022). We show the detailed information about features of the datasets in Table 1.
Metrics and Models. We evaluate the model performance by the following metrics: Overall Accuracy (Acc) measures the proportion of true results (both true positives and true negatives) in the total dataset. Precision (Prec) reflects the ratio of true positive predictions to all positive predictions for each class. Recall (Rec), also known as sensitivity, reflects the ratio of true positive predictions to all actual positives for each class. F-Measure (F1) is the harmonic mean of precision and recall, providing a single score that balances both metrics. We apply the LEKA across a range of models: 1) Tabnet (TN) (Arik and Pfister 2021); 2) TabTransformer (TT) (Huang et al. 2020); 3) Random Forest (RF) (Rigatti 2017); 4) Gradient Boosting Decision Trees (GBDT) (Lin et al. 2023) ; 5) XGBoost (XB) (Chen and Guestrin 2016). We compare the performance outcomes in these tasks both with and without our method.
Baseline Models. We compare the LEKA method with four baseline methods, including: 1) Raw Data: direct using vanilla data; 2) TIFG (Zhang et al. 2024c): feature generation with LLM; 3) KPDDS (Huang et al. 2024): data synthesis with LLM using key points examples; 4) GReaT (Borisov et al. 2022): data synthesis with LLM simulating data subset distributions; 5) FTT (Levin et al. 2022) A TabTransformer pretrained on the source domain and then fine-tuned on the target domain; 6) TTab (Wang and Sun 2022) A transferable tabular Transformer capable of learning from multiple tabular datasets.
Implementation Details. In our setup for data synthesis and model training, we utilize GPT-4o as the query generator, combined with the Exa API to fetch web pages containing datasets from Kaggle and the UCI Machine Learning Repository that may be suitable for knowledge transfer. We extract dataset descriptions from web pages, and GPT-4o assesses their potential for knowledge transfer. Additionally, GPT-4o serves as a generator for executable code, performing up to five code generations. For our models, we configure TN, TT, and FTT with a batch size of $512$ for the VID and TCC datasets and a batch size of $32$ for the BCW dataset, a maximum of $100$ epochs, and employ early stopping with a patience of $20$. The learning rate is set at the default $0.02$ for pytorch_tabnet. For the RF and GBDT models, the number of trees is set to $100$, with GBDT also configured with a learning rate of $0.1$ and a max depth of $3$. TTab is set with a maximum epoch of $50$, a learning rate of $1\times 10^{-3}$, and a weight decay of $1\times 10^{-4}$.
@@
Experiment Results
Overall Performance. The experiment results are shown in Table 3, where we compare our LEKA method in four datasets. In summary:
(1) Compared with baseline models, LEKA significantly improves accuracy, surpassing baseline methods like TIFG and GreAT, with enhancements up to $4.4\%$ in models like GBDT. LEKA consistently achieves top results in precision and recall, indicating its precision in correctly identifying relevant cases while minimizing false positives. The F1 scores under LEKA are notably high, reflecting its effective balance between precision and recall across various models and datasets. These results validate the effectiveness of LEKA ’s retrieval and harmonization strategies and the robustness in diverse application scenarios. Overall, LEKA’s strategic approach to knowledge transfer is particularly advantageous in complex data environments, showcasing its adaptability and efficiency compared to traditional methods.
(2) The LEKA outperforms all baseline methods in almost all metrics and datasets. Specifically, LEKA shows enhancements in accuracy by $2-5\%$ over other methods. It demonstrates notable improvements of up to $4.4\%$ in GBDT for the VID dataset and $5.4\%$ in RF for the BCW dataset. For improvement of overall accuracy, we demonstrate the LEKA’s effective retrieval and harmonization of feature space and marginal probability measures with the target datasets. Meanwhile, the LEKA consistently outperforms baselines in reducing misclassification rates with the highest F1 scores. These results prove that LEKA’s retrieval and data harmonization reduce the likelihood of misclassification by forming a deeper and clearer understanding of the potential relationship between features.
Comparison with transfer learning methods. We then compare our LEKA method with transfer learning methods to demonstrate its effectiveness in enhancing model performance in complex domains. In these scenarios, the transfer learning methods show their reliance on manual data selection and alignment processes. The results demonstrate that LEKA outperforms transfer learning methods FFT and TTab across various metrics and datasets. For example, in the BCW dataset, LEKA improves accuracy by $3.5\%$ and recall by $3.3\%$ compared to the competing methods. This superior performance is consistent across other datasets like VID, HD, and TCC, with notable enhancements in precision and recall, emphasizing LEKA’s effective data harmonization capabilities. This adaptability and efficiency in handling diverse datasets underscore LEKA’s robustness and its significant advantages.
Conclusion
@@
Limitations
We acknowledge the following limitations of our work: (1) The current work is only tested on tabular data. More complex test scenarios are not involved; (2) Despite its automation, LEKA can be computationally intensive. This could limit its applicability in resource-constrained environments or require substantial computational resources to maintain operational efficiency; (3) Adopting the LEKA in tasks with unique requirements may have inherent limitations.

### v2

@@ 5 Experiments
Table 2: Performance comparison of transfer learning methods.
@@
5.1 Experiment Settings
Datasets and Domains.
We evaluate our method primarily on four datasets of medical and economic domains: (1) Breast Cancer Wisconsin (Diagnostic) (BCW) Wolberg et al. (1995), (2) Heart Disease (HD) Janosi et al. (1989), (3) Vehicle Insurance Data (VID) Bhatt (2019), and (4) Telco Customer Churn (TCC) BlastChar (2018). We show the detailed information about features of the datasets in Table 1.
Metrics and Models.
We evaluate the model performance by the following metrics: Overall Accuracy (Acc) measures the proportion of true results (both true positives and true negatives) in the total dataset. Precision (Prec) reflects the ratio of true positive predictions to all positive predictions for each class. Recall (Rec), also known as sensitivity, reflects the ratio of true positive predictions to all actual positives for each class. F-Measure (F1) is the harmonic mean of precision and recall, calculated here as the macro-average. We apply the LEKA across a range of models: 1) Tabnet (TN) Arik and Pfister (2021); 2) TabTransformer (TT) Huang et al. (2020); 3) Random Forest (RF) Rigatti (2017); 4) Gradient Boosting Decision Trees (GBDT) Lin et al. (2023) ; 5) XGBoost (XB) Chen and Guestrin (2016). We compare the performance in these tasks both with and without our method.
Baseline Models.
We compare the LEKA with six baseline methods, including: 1) Raw Data: using vanilla data; 2) TIFG Zhang et al. (2024c): feature generation with LLM; 3) KPDDS Huang et al. (2024): data synthesis with LLM using key point examples; 4) GReaT Borisov et al. (2022): data synthesis with LLM simulating data subset distributions; 5) FTT Levin et al. (2022): A TabTransformer pretrained on the source domain and then fine-tuned on the target domain; 6) TTab Wang and Sun (2022): A transferable tabular Transformer capable of learning from multiple tabular datasets.
Implementation Details.
In our setup for data synthesis and model training, we utilize GPT-4o as the query generator, combined with the Exa API to fetch web pages containing datasets from Kaggle and the UCI Machine Learning Repository that may be suitable for knowledge transfer. We extract dataset descriptions from web pages, and GPT-4o assesses their potential for knowledge transfer. Additionally, GPT-4o serves as a generator for executable code, performing up to five code generations. For our models, we configure TN, TT, and FTT with a batch size of $512$ for the VID and TCC datasets and a batch size of $32$ for the BCW dataset, a maximum of $100$ epochs, and employ early stopping with a patience of $20$. The learning rate is set at the default $0.02$ for pytorch_tabnet. For the RF and GBDT models, the number of trees is set to $100$, with GBDT also configured with a learning rate of $0.1$ and a max depth of $3$. TTab is set with a maximum of $50$ epochs, a learning rate of $1\times 10^{-3}$, and a weight decay of $1\times 10^{-4}$.
@@ 5.1 Experiment Settings
5.2 Experiment Results
Overall Performance.
The experiment results are shown in Table 3, where we compare our LEKA method across four datasets. In summary:
(1) Compared with baseline models, LEKA significantly improves accuracy, surpassing baseline methods like TIFG and GReaT, with enhancements up to $4.4\%$ in models like GBDT. LEKA consistently achieves top results in precision and recall, indicating its precision in correctly identifying relevant cases while minimizing false positives. The F1 scores under LEKA are notably high, reflecting its effective balance between precision and recall across various models and datasets. These results validate the effectiveness of LEKA ’s retrieval and harmonization strategies and its robustness in diverse application scenarios. Overall, LEKA’s strategic approach to knowledge transfer is particularly advantageous in complex data environments, showcasing its adaptability and efficiency compared to traditional methods.
(2) The LEKA outperforms all baseline methods in almost all metrics and datasets. Specifically, LEKA shows enhancements in accuracy by $2-5\%$ over other methods. It demonstrates notable improvements of up to $4.4\%$ in GBDT for the VID dataset and $5.4\%$ in RF for the BCW dataset. For improvement of overall accuracy, we demonstrate the LEKA’s effective retrieval and harmonization of feature space and marginal probability measures with the target datasets. LEKA demonstrates exceptional precision, improving the TT metric for the TCC dataset by over $10\%$ compared to baselines, effectively reducing false positives. Meanwhile, the LEKA consistently outperforms baselines in reducing misclassification rates with the highest F1 scores. These results prove that LEKA’s retrieval and data harmonization reduce the likelihood of misclassification by forming a deeper and clearer understanding of the potential relationship between features.
Comparison with transfer learning methods.
We then compare our LEKA method with transfer learning methods to demonstrate its effectiveness in enhancing model performance in complex domains. In these scenarios, the transfer learning methods show their reliance on manual data selection and alignment processes. The results demonstrate that LEKA outperforms transfer learning methods FFT and TTab across various metrics and datasets. For example, in the BCW dataset, LEKA improves accuracy by $3.5\%$ and recall by $3.3\%$ compared to the competing methods. This superior performance is consistent across other datasets like VID, HD, and TCC, with notable enhancements in precision and recall, emphasizing LEKA’s effective data harmonization capabilities. This adaptability and efficiency in handling diverse datasets underscore LEKA’s robustness and advantages.
6 Conclusion
@@
7 Limitations
We acknowledge the following limitations of our work: (1) The current work has only been tested on tabular data, and more complex test scenarios have not yet been involved; (2) Despite its automation, LEKA can be computationally intensive. This could limit its applicability in resource-constrained environments or require substantial computational resources to maintain operational efficiency; (3) Adopting the LEKA in tasks with unique requirements may have inherent limitations.
Acknowledgements
Dr. Kunpeng Liu is supported by the National Science Foundation (NSF) via the grant numbers 2426339 and 2348485. Dr. Yanjie Fu is supported by the National Science Foundation (NSF) via the grant numbers 2426340, 2416727, 2421864, 2421865, 2421803, and National Academy of Engineering Grainger Foundation Frontiers of Engineering Grants.